Solve the given differential equation by using an appropriate substitution. The DE is of the form dy dx = f(Ax + By + C)



dy/dx = (x + y + 1)^2

1
Expert's answer
2021-11-10T17:07:49-0500

"\frac{dy}{dx}=(x+y+1)²"

Let (x+y+1) = t

Differentiating with respect to x ,

1+ "\frac{dy}{dx} = \frac{dt}{dx}"

=> "\frac{dy}{dx} = \frac{dt}{dx}-1"

So the differential equation becomes

"\frac{dt}{dx}-1 = t²"

=> "\frac{dt}{dx} = (1+t²)"

=> "\frac{dt}{1+t²} = dx"

Integrating ,

"\int \frac{dt}{1+t²} = \int dx"

=> "tan^{-1}t = x +" C , where C is integration constant.

=> tan"^{-1}" ( x + y + 1 ) = x + C

So the solution of the given differential equation is tan"^{-1}" (x + y + 1) = x + C where C is integration constant.